OpenStudy (anonymous):

A spinner, with eight equal sections numbered 1 through 8, is spun 50 times. The number 3 is recorded 8 times. Determine the experimental probability of spinning a 3.

OpenStudy (anonymous):

8/50 = 4/25 = 0.16

OpenStudy (anonymous):

Thank you.